# Monocular pose estimation of articulated surgical instruments in open surgery

### v1

@@ 2.1 Overview of the Proposed Pipeline
The proposed pipeline addresses the unique challenges of the surgical environment, such as the scarcity of annotated real-world data, the geometry of articulated instruments, and the presence of occlusions.
The core goal is to enable accurate pose estimation without extensive manual annotation of real surgical data. To achieve this, a three-component pipeline is proposed, consisting of synthetic data generation, a tailored pose estimation framework, and a training strategy that leverages both synthetic and real unannotated data.
@@ 2.2 Synthetic Data Generation
We decided to create our own CAD models of the surgical instruments instead of relying on existing CAD models. This decision was driven by the need for highly realistic and accurate representations, as well as the unavailability of CAD models.
To prepare surgical instruments for photogrammetry, we spray them with a dust spray to minimize glare and reflections that could interfere with the scanning process. We found that dry shampoo worked better as a dust spray than specialized sprays for 3D scanning. It resulted in a better 3D reconstruction and is more cost-effective. The tools are then placed on a spinning table and photographed 360 degrees at three different heights. This ensures that every detail and surface of the instruments is accurately captured.
The captured images are processed using Reality Capture [14], a professional photogrammetry software. Once the initial 3D meshes are generated, we perform a series of refinement steps to enhance their quality and realism. We use Blender [8], a popular free and open source 3D modeling and animation software, to clean the meshes by removing noise and artifacts introduced during the photogrammetry process. Additionally, we employ Blender’s smoothing and hole-filling tools to improve the mesh surface quality.
@@ 2.2.2 Hand-Object Interaction Modeling
Occlusions of hand grasps are commonly encountered in surgical scenarios. We simulate them with a generative hand-grasp model called ContactGen [36] in combination with the MANO hand mesh [48]. We used the model from the official implementation, which was trained on the GRAB dataset [54, 11]. The GRAB dataset includes real human grasps for 51 objects from 10 different subjects.
@@ 2.2.3 Rendering Pipeline
In total, we generate 80,000 synthetic images using our rendering pipeline, utilizing BlenderProc [16]. Each image is accompanied by precise ground-truth annotations, including the 6D pose of the surgical instruments and the articulation angles. This comprehensive synthetic dataset serves as a valuable resource for training our pose estimation network. A sample from this dataset is illustrated in Figure 5.
@@ 2.3 Pose Estimation Framework
Our pose estimation framework follows a two-stage approach, consisting of an object detection stage followed by a pose estimation stage. The first stage involves detecting the objects of interest in the input image. The second stage takes the detected object crops and estimates the 6D pose, object class, and articulation angle.
The object detection stage is handled by YOLOv8 [27], which is trained to detect and localize the surgical tools in the input image. The detector provides bounding boxes around the detected tools, along with confidence scores indicating the likelihood of each surgical tool class.
The pose estimation stage takes as input cropped object regions from the detector and processes them through the pose estimation network. The pose estimation network estimates the 6D pose, object class, and articulation angle of the surgical tool.
@@ 2.3.1 Direct 6D Object Pose Estimation
As briefly mentioned in Related Work, our pose estimation method draws inspiration from GDR-Net [60] for direct 6D object pose estimation, specifically the GDRNPP implementation [37], which incorporates several improvements over the original GDR-Net presented in the conference version. It is worth noting that GDRNPP won most awards in the 2022 BOP challenge [52] for “6D localization of seen objects” and “2D detection of seen objects”, demonstrating its state-of-the-art performance. We adopt several key components from GDR-Net, including the 3D rotation parameterization, 3D translation parameterization, 2D-3D dense correspondence maps, Patch-PnP, and the multitask 6D pose loss, which are described in the following paragraphs.
@@
For the 3D rotation, we adopt the 6D rotation representation [67] instead of commonly used representations like quaternions or Euler angles. Quaternions and Euler angles suffer from ambiguities and discontinuities, which can hinder the learning process and lead to suboptimal pose estimates. The 6D rotation representation provides a continuous and unambiguous way to represent rotations.
The 6D rotation representation consists of two 3D vectors, denoted $r_{1}$ and $r_{2}$, which correspond to the first two columns of the rotation matrix $\mathbf{R}{\in}\textit{SO}(3)$. The network predicts these two columns using the Patch-PnP module. To ensure the orthogonality and validity of the rotation matrix, Gram-Schmidt process is applied to the predicted columns. This process orthogonalizes the columns and computes the third column to complete the rotation matrix. The resulting rotation matrix $R=[r_{1},r_{2},r_{3}]$ is guaranteed to be orthogonal and represents a valid 3D rotation.
In the context of working with object crops, we focus on predicting the allocentric representation of an object’s rotation [29]. This approach is independent of the viewer’s perspective and assumes that the camera has rotated from its original position to point toward the center of the object’s Region of Interest (RoI). This representation eliminates the need to account for variable camera-object positioning, simplifying the rotational estimation process. Once the allocentric rotation is determined, it can be converted into the true 3D rotation (egocentric representation) by using the object’s 3D translation and the camera’s intrinsic parameters $K$ [29].
@@
Direct prediction of translation from a Region of Interest (RoI) is impossible without additional information. To address this, we use a parametrization technique known as the scale-invariant translation estimate (SITE) [32]. SITE represents the translation as a combination of the 2D center coordinates of the object in the image plane and the depth of the object relative to the camera. The network predicts three components: $(\delta_{x},\delta_{y},\delta_{z})$.
@@
The network generates a dense 2D-3D correspondence map [32, 24], which is an intermediate geometric feature that establishes a relationship between each object pixel in the 2D image and its corresponding 3D coordinate on the object model. These 3D coordinates are normalized using the dimensions of a tight 3D bounding box that encloses the mesh. Utilizing this correspondence map, we can apply PnP-RANSAC algorithms to find a plausible 6D pose of the object.
@@ 2.3.3 Pose Model Architecture
Fig. 6: Pose Model Architecture: The input image is cropped at the Region of Interest (RoI) and processed by the Multitask Feature Regression. The output, $M_{2D-3D}$, is then forwarded to Patch-PnP during training or to PnP-RANSAC during testing.
The intermediate feature regression stage is based on the Feature Pyramid Network (FPN) [33] with a Vision Transformer (ViT) backbone pretrained following the method described in [43] on the LVD-142M dataset. The features extracted by the ViT backbone are fed into a Pyramid Pooling Module (PPM) [66] before being fed back to the FPN. The object category head is attached to the PPM output for early branching [19]. The fused feature map from the FPN is directed to three heads: articulation, mask, and 2D-3D Dense Correspondence ($M_{2D-3D}$).
The pose regression stage varies depending on the phase. During training, the $M_{2D-3D}$ encoding is processed by Patch-PnP [60], a differentiable pose estimator. During testing, the $M_{2D-3D}$ encoding is converted to 2D-3D pairs, and the pose is estimated using a PnP-RANSAC variant. The final output consists of the pose parameters ($R_{3x3},t_{3}$), object class, and articulation angle.
@@ 2.3.4 Training Procedure
The training process involves the following steps:
  1. 1.
    Train the pose estimation network and the object detection network independently on synthetic data.
  2. 2.
    Create pseudo-labels for the object detector using the pose estimation network and retrain the object detector on the pseudo-labels.
  3. 3.
    Create pseudo-labels for the pose estimator using the updated object detector and retrain the pose estimator on the pseudo-labels.
  4. 4.
    Repeat steps 2-3 for multiple iterations to progressively refine the networks’ performance on real-world data.
The details of each step are discussed in the following subsections.
Stage 1: Training on Synthetic Data
In the first stage, we train the pose estimation network and the object detection network independently on the synthetic dataset generated using the pipeline described in Section 2.2.
@@
    We apply object tracking on the real video data using the BoT-SORT [2] tracking algorithm.
@@
    We select frames with high detection confidence and at least three consecutive tracking detections as pseudo-labels for object detection.
@@
    The selected frames are passed to the pose estimator, which estimates the object pose, class, and articulation angle. The resulting mesh and pose are reprojected onto the image plane. The resulting reprojected bounding box is then used as the new pseudo-label for the object detection network. This process is illustrated in Figure 7.
@@
    Retrain the object detector on the pseudo-labeled data and a fraction of synthetic data.
@@
    Using the refined object detection model, we obtain image crops of the detected objects from the real video data.
@@
    We run the pose estimation network on these image crops to predict the 6D pose, object class, and articulation angle.
@@
    We filter the pose predictions based on the following criteria:
    • 3..1
      High class confidence
    • 3..2
      Low PnP-RANSAC outlier count
    • 3..3
      Low reprojection error
    The pose predictions that satisfy these criteria are selected as pseudo-labels for pose estimation.
@@
    Retrain the pose estimator on the pseudo-labeled data and a fraction of synthetic data.
@@ 3.1 Experimental Setup
Our experiments are implemented with the PyTorch deep learning framework. We train our models end-to-end using the Ranger optimizer [35, 65, 63], which combines Rectified Adam (RAdam) with Lookahead. A batch size of 120 is used during training, along with a base learning rate of 1e-5. We employ a cosine annealing learning rate schedule [38], which gradually reduces the learning rate after approximately 72% of the training iterations.
During training, we apply various data augmentation techniques with a base probability of 0.8. These augmentations are taken from the GDRNPP implementation [60, 37] and include: Gaussian blur, Sharpness enhancement, Contrast enhancement ,Brightness adjustment ,Color enhancement, Addition of random values to pixel intensities, Channel-wise inversion, Multiplication of pixel intensities ,Addition of Gaussian noise, Linear contrast adjustment, and Grayscale conversion.
For inference, we use the Progressive-X [6] PnP-RANSAC scheme, which utilizes Graph-Cut RANSAC [5].
@@ 3.1.2 Evaluation Metrics
Due to the lack of ground truth 6D pose annotations in our real-world dataset, we employ a 2D reprojection-based evaluation method. This approach compares the reprojected mask of the object, obtained using the estimated pose, class, and articulation angle, with the manually annotated segmentation masks. However, directly comparing the full reprojection masks to the annotated tool masks may not provide an accurate assessment of the pose estimation accuracy, as the annotated tool masks may be partially occluded by the surgeon’s hands. To account for these occlusions, we incorporate the hand masks into our evaluation process.
@@
    We use the Average Precision ($AP$) metric for segmentation following the method used in the COCO 2020 Object Segmentation Challenge [34] and BOP 2022 Challenge [52]. For each object, we calculate a per-object Average Precision ($AP_{O}$) by averaging the $AP$ across the following Intersection over Union (IoU) thresholds: [0.5, 0.55, 0.6, 0.65, 0.7, 0.75, 0.8, 0.85, 0.9, 0.95]. The IoU is evaluated between the manually subtracted annotated tool mask and the manually subtracted reprojected tool mask. The overall performance metric ($AP$) is then computed by averaging the $AP_{O}$ scores for all objects.

### v2

@@ 3.1 Overview of the Proposed Pipeline
The proposed framework addresses the unique challenges of the surgical environment, such as the scarcity of annotated real-world data, the geometry of articulated instruments, and the presence of occlusions.
The core goal is to enable accurate pose estimation without extensive manual annotation of real surgical data. To achieve this, a three-component pipeline is proposed, consisting of synthetic data generation, a tailored pose estimation network, and a training strategy that leverages both synthetic and real unannotated data.
@@ 3.2 Synthetic Data Generation
We developed custom CAD models for our surgical instruments due to a lack of suitable existing options and the need for highly realistic and accurate tool representations. To achieve this level of detail and accuracy, we employed photogrammetry techniques to capture the intricate features of the actual instruments.
To prepare the surgical instruments for photogrammetry, we spray them with a dust spray to minimize glare and reflections that could interfere with the scanning process. Dry shampoo, surprisingly, outperformed specialized 3D scanning sprays, providing better reconstructions at a lower cost. The tools are then placed on a spinning table and photographed 360 degrees at three different heights. This approach ensures that every detail and surface of the instruments is accurately captured, providing us with the high-quality data needed to create our custom CAD models.
The captured images are processed using Reality Capture [RealityCapture], a professional photogrammetry software. Once the initial 3D meshes are generated, we perform a series of refinement steps to enhance their quality and realism. We use Blender [BlenderFoundation2023], a popular free and open source 3D modeling and animation software, to clean the meshes by removing noise and artifacts introduced during the photogrammetry process. Additionally, we employ Blender’s smoothing and hole-filling tools to improve the mesh surface quality.
@@ 3.2.2 Hand-Object Interaction Modeling
Occlusions of hand grasps are commonly encountered in surgical scenarios. We simulate them with ContactGen [liu2023contactgen], a generative hand-grasp model, in combination with the MANO hand mesh [MANO:SIGGRAPHASIA:2017]. We used the model from the official implementation, which was trained on the GRAB dataset [GRAB:2020, Brahmbhatt_2019_CVPR]. The GRAB dataset includes real human grasps for 51 objects from 10 different subjects.
@@ 3.2.3 Rendering Pipeline
In total, we generate 80,000 synthetic images using our rendering pipeline, utilizing BlenderProc [Denninger2023]. Each image is accompanied by precise ground-truth annotations, including the 6D pose of the surgical instruments and the articulation angles. This comprehensive synthetic dataset serves as a valuable resource for training our pose estimation network. A sample from this dataset is illustrated in Figure 5.
@@ 3.3 Pose Estimation Framework
Our pose estimation framework combines object detection and 6D pose estimation stages with a domain adaptation strategy designed to bridge the gap between synthetic and real-world surgical data. This approach combines established, state-of-the-art components with novel adaptations tailored for the unique challenges of surgical tool pose estimation.
For the initial object detection stage, we employ YOLOv8 [Jocher_Ultralytics_YOLO_2023], a state-of-the-art object detection framework recognized for its speed and accuracy. We adapt its standard training process to incorporate our synthetic-to-real domain adaptation strategy, which is detailed further in Section 3.3.4.
The subsequent pose estimation stage receives cropped object regions identified by the detector. It then processes them using our pose estimation network. This network builds upon the strong foundation of GDR-Net [wang2021gdr] and its enhanced implementation, GDRNPP [liu2022gdrnpp_bop], known for state-of-the-art performance in general 6D object pose estimation. From these baseline models, we adopt several key components proven effective for pose estimation: we utilize the continuous 6D rotation parameterization proposed by [zhou2019continuity] for stable orientation representation, incorporate the differentiable Patch-PnP module [wang2021gdr] to enable end-to-end pose regression during training, and employ multi-task loss functions for comprehensive optimization of pose-related objectives.
For domain adaptation, we integrate pseudo-labeling cycles inspired by iterative refinement strategies in prior work (e.g., [hein2021towards]). This iterative process is employed during training to progressively refine model predictions on unlabeled real-world data, thereby mitigating the performance gap caused by the domain shift from the annotated synthetic source data.
Our novel contributions include architectural and methodological adaptations. Architecturally, we introduce a multitask feature regression module (Figure 6) for simultaneous pose, category, and articulation angle estimation from image features. We also incorporate mechanisms to handle tool articulation and type variability, enabling generalization to new tools without requiring specific geometric models. Methodologically, our domain adaptation strategy extends pseudo-labeling techniques to simultaneously improve detection and pose estimation for articulated tools, using only synthetic data as the annotated source.
@@ 3.3.1 Direct 6D Object Pose Estimation
As briefly mentioned in Related Work, our pose estimation method draws inspiration from GDR-Net [wang2021gdr] for direct 6D object pose estimation, specifically the GDRNPP implementation [liu2022gdrnpp_bop], which incorporates several improvements over the original GDR-Net presented in the conference version. It is worth noting that GDRNPP won most awards in the 2022 BOP challenge [sundermeyer2023bop] for “6D localization of seen objects” and “2D detection of seen objects”, demonstrating its state-of-the-art performance. We adopt several key components from GDR-Net, including the 3D rotation parameterization, 3D translation parameterization, 2D-3D dense correspondence maps, Patch-PnP, and the multitask 6D pose loss, which are described in the following paragraphs.
@@
For the 3D rotation, we adopt the 6D rotation representation [zhou2019continuity] instead of commonly used representations, such as quaternions or Euler angles. Quaternions and Euler angles suffer from ambiguities and discontinuities, which can hinder the learning process and lead to suboptimal pose estimates. The 6D rotation representation provides a continuous and unambiguous way of representing rotations.
Fig. 6: Pose Estimation Model Framework: An input RGB image is first cropped to the Region of Interest (RoI) and processed by the Multitask Feature Regression Network, which uses a Vision Transformer (ViT) backbone for feature extraction. The extracted features pass through a UPerNet architecture with a Pyramid Pooling Module (PPM) to aggregate multiscale contextual information, then branch into four output heads: object category classification, articulation angle regression, object masks, and 2D-3D dense correspondence mapping. The object masks help stabilize training and facilitate domain adaptation. The resulting dense correspondence output ($M_{2D-3D}$) is concatenated with CoordConv layers and then sent either to the Patch-PnP module (during training) or to PnP-RANSAC (during inference) for pose estimation.
The 6D rotation representation consists of two 3D vectors, denoted $r_{1}$ and $r_{2}$, which correspond to the first two columns of the rotation matrix $\mathbf{R}{\in}\textit{SO}(3)$. The network predicts these two columns using the Patch-PnP module. To ensure the orthogonality and validity of the rotation matrix, the Gram-Schmidt process is applied to the predicted columns. This process orthogonalizes the columns and computes the third column to complete the rotation matrix. The resulting rotation matrix $R=[r_{1},r_{2},r_{3}]$ is guaranteed to be orthogonal and represents a valid 3D rotation.
In the context of working with object crops, we focus on predicting the allocentric representation of an object’s rotation [kundu20183d]. This approach is independent of the viewer’s perspective and assumes that the camera has rotated from its original position to point toward the center of the object’s Region of Interest (RoI). This representation eliminates the need to account for variable camera-object positioning, simplifying the rotational estimation process. Once the allocentric rotation is determined, it can be converted into the true 3D rotation (egocentric representation) by using the object’s 3D translation and the camera’s intrinsic parameters $K$ [kundu20183d].
@@
Direct prediction of translation from a Region of Interest (RoI) is impossible without additional information. To address this, we use a parameterization technique known as the scale-invariant translation estimate (SITE) [li2019cdpn]. SITE represents the translation as a combination of the 2D center coordinates of the object in the image plane and the depth of the object relative to the camera. The network predicts three components: $(\delta_{x},\delta_{y},\delta_{z})$.
@@
The network generates a dense 2D-3D correspondence map [li2019cdpn, hodan2020epos], which is an intermediate geometric feature that establishes a relationship between each object pixel in the 2D image and its corresponding 3D coordinate on the object model. These 3D coordinates are normalized using the dimensions of a tight bounding box enclosing the object’s mesh. This normalized coordinate representation is referred to as NOCS maps [wang2019normalized]. Generating these 2D–3D correspondences for articulated objects can be done in two ways: (1) fixing correspondences per articulation angle via its own bounding box, or (2) normalizing by the bounding box of the fully articulated configuration. We adopt the second approach to avoid maintaining a lookup table. We then use CoordConv layers [liu2018intriguing] by concatenating (x, y) coordinates to the dense correspondences as channels, providing explicit 2D positional encoding. These layers are particularly useful after cropping the images. The dense correspondence map can then be used with PnP-RANSAC to solve for the object’s 6D pose.
@@ 3.3.3 Pose Model Architecture
The intermediate feature regression stage is based on UPerNet [xiao2018unified] with a Vision Transformer (ViT) backbone pretrained following the method described in [oquab2023dinov2] on the LVD-142M dataset. The features extracted by the ViT backbone are fed into a Pyramid Pooling Module (PPM) [zhao2017pyramid] before being fed back to the FPN. The object category head is attached to the PPM output for early branching [elhoseiny2015convolutional]. The fused feature map from the FPN is directed to three heads: articulation, mask, and 2D-3D Dense Correspondence ($M_{2D-3D}$).
The pose regression stage varies depending on the phase. During training, the $M_{2D-3D}$ encoding is processed by Patch-PnP [wang2021gdr], a differentiable pose estimator. During testing, the $M_{2D-3D}$ encoding is converted to 2D-3D pairs, and the pose is estimated using a PnP-RANSAC variant. The final output consists of the pose parameters ($R_{3x3},t_{3}$), object class, and articulation angle.
@@ 3.3.4 Training Procedure
The training involves iterative domain adaptation, starting with independent training on synthetic data. Pseudo-labels generated from one network’s predictions are used to retrain the other, and this process is repeated twice. Details are provided in the following subsections.
Step 1: Training on Synthetic Data
In the first step, the pose estimation network and the object detection network are trained independently on the synthetic dataset generated using the pipeline described in Section 3.2.
@@
    Object tracking is applied to the real video data using the BoT-SORT tracking algorithm
@@
    Frames with high detection confidence and at least three consecutive tracking detections are selected as pseudo-labels for object detection.
@@
    The selected frames are processed by the pose estimator to predict the object pose, class, and articulation angle. The resulting mesh and pose are reprojected onto the image plane, and the reprojected bounding box is used as a new pseudo-label for the object detection network. This process is illustrated in Figure 7.
@@
    The object detector is retrained on the pseudo-labeled data along with a fraction of the synthetic data.
@@
    Image crops of the detected objects are obtained from the real video data using the refined object detection model.
@@
    The pose estimation network is applied to these image crops to predict the 6D pose, object class, and articulation angle.
@@
    Pose predictions are filtered based on class confidence, PnP-RANSAC outlier count, and reprojection error. Pose predictions that satisfy these filtering criteria are selected as pseudo-labels for retraining the pose estimation network. The specific numerical thresholds used for filtering are detailed in Section 4.1.2. and a broader discussion of these criteria, along with potential future work, is provided in the Discussion section.
@@
    The pose estimator is retrained on the pseudo-labeled data along with a fraction of synthetic data.
Fig. 8: Annotated segmentation masks of surgical tools and hands.
@@ 4.1 Experimental Setup
Our experiments are implemented with the PyTorch deep learning framework. We train our models end-to-end using the Ranger optimizer [liu2019variance, zhang2019lookahead, yong2020gradient], which combines Rectified Adam (RAdam) with Lookahead. A batch size of 120 is used during training, along with a base learning rate of 1e-5. We employ a cosine annealing learning rate schedule [loshchilov2016sgdr], which gradually reduces the learning rate after approximately 72% of the training iterations.
During training, we apply various data augmentation techniques with a base probability of 0.8. These augmentations are taken from the GDRNPP implementation [wang2021gdr, liu2022gdrnpp_bop] and include: Gaussian blur, sharpness enhancement, contrast enhancement, brightness adjustment, color enhancement, addition of random values to pixel intensities, channel-wise inversion, multiplication of pixel intensities, addition of Gaussian noise, linear contrast adjustment, and grayscale conversion.
For inference, we use the Progressive-X [barath2019progressive] PnP-RANSAC scheme, which utilizes Graph-Cut RANSAC [barath2018graph].
@@ 4.1.1 Dataset and Labels
4.1.2 Thresholds for Filtering Predictions
The thresholds used for filtering pose predictions during the domain adaptation step (Section 3.3.4, Step 3) were chosen after evaluating their impact on the accuracy of pseudo-label generation during synthetic-to-real adaptation. The criteria are:
  • 1.
    Class confidence: Predictions are retained if their class confidence score exceeds 0.999.
  • 2.
    PnP-RANSAC outlier count: Only predictions with an outlier count below 30 are accepted.
  • 3.
    Reprojection error: Predictions with a Jaccard index above 0.3 are selected.
@@ 4.1.3 Evaluation Metrics
Due to the lack of ground truth 6D pose annotations in our real-world dataset, we employ a 2D reprojection-based evaluation method [brachmann2016uncertainty]. This approach compares the reprojected mask of the object, obtained using the estimated pose, class, and articulation angle, with the manually annotated segmentation masks. However, directly comparing the full reprojection masks to the annotated tool masks may not provide an accurate assessment of the pose estimation accuracy, as the annotated tool masks may be partially occluded by the surgeon’s hands. To account for these occlusions, we incorporate the hand masks into our evaluation process.
@@
    We use the Average Precision ($AP$) metric for segmentation following the method used in the COCO 2020 Object Segmentation Challenge [lin2014microsoft] and BOP 2022 Challenge [sundermeyer2023bop]. For each object, we calculate a per-object Average Precision ($AP_{O}$) by averaging the $AP$ across the following Intersection over Union (IoU) thresholds: [0.5, 0.55, 0.6, 0.65, 0.7, 0.75, 0.8, 0.85, 0.9, 0.95]. The IoU is evaluated between the manually subtracted annotated tool mask and the manually subtracted reprojected tool mask. The overall performance metric ($AP$) is then computed by averaging the $AP_{O}$ scores for all objects.